# ARN: Analogical Reasoning on Narratives

### v3

@@ 1 Introduction
Analogical reasoning is a core cognitive skill unique to humans (Penn et al., 2008; Hofstadter, 2001), defined as the ability to perceive and utilize the similarities between situations or events based on (systems of) relations rather than surface similarities (Holyoak, 2012; Gentner et al., ). The dichotomy between relational and surface similarity is illustrated in Figure 1: while the narratives $Q$ and $N$ overlap in terms of characters, locations, and actions (surface similarity), they fail to create a system of relational correspondences and thus are disanalogous as shown by their high-level message. Meanwhile, $Q$ and $A$ are dissimilar on the surface, while forming a coherent relational system of correspondences through the high-level message: no pain, no gain. Mappings based on systems of relations (system mappings) have priority over independent relations (Gentner et al., 1993). Further, depending on whether the corresponding situations/events share the same domain or not, analogies and disanalogies can be qualified as near or far (Gentner, 1983).
@@
Analogy enables creative inferences, explanations, and generalization of knowledge and has been used for scientific inventions (Dunbar and Klahr, 2012), solving problems (Gick and Holyoak, 1980), and policy-making (Houghton, 1998). It has also been the subject of cognitive theories and studies about humans for common processes, such as retrieval of memories (Wharton et al., 1994) and problem-solving (Gick and Holyoak, 1980), mostly leveraging narratives as their experimental medium (e.g. Gentner and Toupin, 1986; Gentner et al., 1993; Wharton et al., 1994), given their multi-tiered nature and potential for abstraction. However, these studies focus on humans or models expecting well-structured inputs (Lu et al., 2022), are small-scale (Ichien et al., 2020, Table 5), and involve manually curated narratives, lacking an indication of how they can be applied to create scalable benchmarks for evaluating computational models. Nevertheless, borrowing from narratology theories Mani (2013); Gardner (2003), elements of narratives may enable analogical theories to be scaled and formalized as computational frameworks and corresponding benchmarks; however, this avenue has not been explored to date.
@@
Meanwhile, analogical reasoning has also been relatively popular in natural language processing (NLP), typically framed as an intelligence test for models compared against humans. So-called word-based, proportional analogies of the form $(A:B::C:D)$ (e.g. Mikolov et al., 2013a, b; Gladkova et al., 2016; Ushio et al., 2021) are often used to measure the potential of word embeddings and language models in terms of analogical reasoning. Recent studies (Webb et al., 2023) show a strong ability of state-of-the-art (SOTA) large language models (LLMs) to discover proportional word analogies, though this skill degrades with higher complexity (Wijesiriwardene et al., 2023) or associative phrasing of the input (Stevenson et al., 2023). Shifting toward more complex settings, narrative-based analogy benchmarks that involve system mappings rather than simple word-based relational mappings, and are aligned with cognitive theories have been rarely considered (Nagarajah et al., 2022; Wijesiriwardene et al., 2023), with limitations in scope and generalizability. Thus, we note a significant gap between the expressivity of cognitive theories of analogies and the present analogical reasoning benchmarks in NLP. Specifically, existing computational frameworks seldom include analogy over narratives and do not provide a formalization of key concepts such as system mappings and near/far analogs. Consequently, the benchmarks used to evaluate SOTA models do not provide the ability for systematic exploration of the sensitivity of model performance to different analogical categories beyond word-level relations.
@@ 2 Related Work
Many categorizations exist for analogies, with a common distinction between surface and relational (system) mappings (e.g. Halford, 1992; Gentner and Rattermann, 1991; Gentner, 1982; Premack, 1983). Gentner (1983)’s Structural Mapping Engine (SME) defines domains and situations as systems of objects, object attributes, and relations between objects, and distinguishes between mappings formed based on each. SME defines analogies as mappings based on relations rather than objects (attributes), prioritizing interconnected systems of mappings across domains (system mappings). Similarly, Holyoak and Thagard (1996) distinguish between three mapping categories of increasing complexity: attribute, relational, and system. Each mapping category can be formalized using a propositional representation with predicates (e.g., Mentor of) and filler subjects (e.g., Kim, Emily). Aligned with both frameworks, we distinguish between two categories of surface mappings covering all categories of lower-order mappings (e.g., based on objects and object attributes), and system of relational mappings (system mappings) that create a coherent system of correspondences across two situations.
System mappings form analogies, having a higher priority over other mappings (Gentner, 1983; Holyoak and Thagard, 1996); yet, they can co-exist with surface mappings. Surface mappings decrease the cognitive load for recognizing system mappings since they create within-domain (near) analogies, also known as literal similarity (Gentner, 1983). Cross-domain (far) analogies, not forming surface mappings, are more challenging to identify (Alexieva and Hristova, 2017; Green et al., 2006, 2008, 2010), promote relational thinking (Vendetti et al., 2014), and are seen as more sound when making arguments (Gentner et al., 1993). Conversely, the sole existence of surface mappings yields (mere appearance) near disanalogies, while the absence of both surface and system mappings is a far disanalogy, or dissimilarity (Gentner, 1983). While the distinction between near and far (dis)analogies is common in popular cognitive psychology (CogPsy) frameworks, it is unclear how to operationalize them at scale for narratives, a gap we bridge in this paper.
@@
To understand what may form surface and system mappings in narratives, we review frameworks that describe narratives through their key elements. Bal (2017) distills four narrative elements: actors, events, time, and location. Mani (2013) distinguish between characters, time, and plot. Building upon these categorizations, Vossen et al. (2021) expands the scope to include sequentiality and focalizer in addition to events and characters. More comprehensive and more widely adopted, Gardner (2003) consider characters, plot, theme, setting, point of view, and style. To our knowledge, no prior work has leveraged such elements to infer surface and system similarities across narratives. We base our categorization of elements on Gardner (2003), given its comprehensiveness, to form surface and system mappings, inferring analogies and disanalogies across narratives.
@@
For both models and humans, evaluating analogical reasoning as an intelligence test (Hesse, 1965; Mitchell, 2021) is often posed as a word-based proportional task ($A:B::C:D$); having $Happy:Sad::C:D$, and asking to infer the missing $Angry:Calm$ or $Calm$ given $Happy:Sad::Angry:D$. Proportional analogy benchmarks (e.g. Turney et al., 2003; Jurgens et al., 2012; Mikolov et al., 2013a, b, c; Gao et al., 2014; Gladkova et al., 2016; Czinczoll et al., 2022; Kotchian and Simmons, 2012; Tutors, ) differ in scale (up to thousands of samples), number of relations (between one and few dozens), and domains (general domain or science). Depending on the context, SOTA LLMs perform worse than humans (Collier et al., 2022; Mitchell et al., 2023), or outperform humans (which might be artifacts of training data; Lewis and Mitchell, 2024) in word-based analogy tasks; however, they fail to perform well in story analogies (Webb et al., 2023). Notably, while cognitive theories of analogy often focus on narratives and study the distinction between surface, relational, and system similarities, existing work in NLP has a narrower focus as far as analogies between word pairs. In this study, we extend the focus from word-based relational analogies to relational system analogies in narratives.
@@
Meanwhile, focusing on the limited studies of matching, retrieving, or completing more complex analogies in narratives in CogPsy (Ichien et al., 2020): Wharton et al. (1994) develop 14 sets of four stories with different themes and events, used to compare human analogies between themes and events; Wijesiriwardene et al. (2023) study analogies in stories using negation, entailment, and explanations, using available datasets for each; Gentner and Toupin (1986) study the effect of systematicity and surface similarity of narratives’ characters on analogical reasoning with 54 narratives; Gick and Holyoak (1980) study analogical reasoning in problem-solving and assess the effect of similarities in problems and solutions separately. These tasks consider narrative analogies formed by humans, are thus limited in size (dozens of samples), and do not exploit the richness of the frameworks on narrative elements. Although large-scale story corpora exist (e.g. Mostafazadeh et al., 2016; Storks et al., 2021; Andrus et al., 2022), analogies are extremely scarce in free texts (Sultan and Shahaf, 2022), and these resources are not adequate for evaluating analogical reasoning (Wijesiriwardene et al., 2023) since they do not consider the theoretical dimensions of analogical reasoning from a CogPsy perspective, which is a gap we cover in this study. Most similar to our work, Nagarajah et al. (2022) formalize six cognitively-inspired dimensions of analogical reasoning over hundreds of Aesop’s fables. However, while these dimensions can be mapped to our framework, their fine granularity limits their application to the set of fables they are devised for. Instead, we formulate a comprehensive framework that formalizes analogies and disanalogies by using narrative elements from narratology. This framework explicitly connects to the CogPsy theoretical dimensions of analogical reasoning frameworks, resulting in the inception of the first extensive, theory-grounded benchmark for analogical reasoning within narratives. Concurrently to our work, Jiayang et al. (2023) focus on analogies in narratives by providing a large-scale benchmark built on top of existing resources; however, they focus specifically on the distinction between surface and relational similarities. We go one step further by covering system mappings and the possible interactions between near/far analogies and distractors to paint a more accurate picture of models’ analogical reasoning abilities.
@@ 3 $ARN$ Framework
Inspired by analogical reasoning frameworks in CogPsy (e.g. Wharton et al., 1994; Gentner et al., 1993), we formulate analogical reasoning over narratives as a task where input consists of a query narrative $Q$ and two candidate narratives, $A$ and $N$ (see Figure 1). The goal is to select the narrative $A$ that is analogous to $Q$ rather than the disanalogous one $N$, based on system mappings. We frame it as a binary classification task to enable an unambiguous evaluation of the analogical reasoning abilities of models and humans (Frank, 2023). We realize this task through a compositional framework with three components (Figure 2): 1. extraction of elements of narratives, such as characters, actions, and high-level messages (see Section 3.1); 2. formation of surface and system mappings based on those elements (see Section 3.2); 3. task formulation, by defining (dis)analogies based on the mappings and partitioning the data points according to their semantic distance to the query narratives (see Section 3.3).
@@ 3.1 Elements of Narratives
We take the proverbial or high-level message of a narrative as an instantiation of its overarching theme (Gardner, 2003), which encapsulates individual components of narratives into a coherent idea that the writer wants to convey. More formally, proverbs are short sentences containing wisdom, truth, morals, and traditional views in a metaphorical form (Mieder, 1993). For instance, query narrative $Q$ in Figure 1 is associated with the proverb: no pain, no gain.
@@ 3.2 Mappings between Narratives
Unlike surface mappings that focus on similarities between independent elements, system mappings are formed by interconnected correspondences of relations across two narratives (Gentner, 1983). This system of correspondences can be characterized and formed by leveraging narratives’ proverbs (Wijesiriwardene et al., 2023); hence, two narratives with the same proverb are connected by system mappings. Similar to the moral dimension in Nagarajah et al. (2022), proverbs provide an abstract summary of the causal structure of a narrative, acting as mediators for the transition between two analogs.
@@ 3.4 Generalizability of ARN to Narrative Analogies
The purpose of $ARN$ is to provide a generalizable framework to study analogical reasoning in narratives grounded in both narratology and analogical reasoning in CogPsy research. The provided set of elements is universal in narratives (Gardner, 2003): entities’ attributes or surface relations, proverbs as overarching messages of narratives, together with their corresponding surface and system mappings, can be broadly found in any narrative. This universality of its components makes $ARN$ an adequate framework to describe other existing datasets. To illustrate this point, we describe how two existing benchmarks for narrative analogies can be mapped to the components of $ARN$. StoryAnalogy (Jiayang et al., 2023) uses characters, actions, and relationships of narratives to represent and distinguish between entity and relational similarities while not covering higher-order mappings of analogical messages. The entity and relational elements of StoryAnalogy directly correspond to the elements in $ARN$, whereas higher-order elements and mappings, such as proverbs, are not present in this dataset. Similarly, Nagarajah et al. (2022) capture structural similarities between Aesop’s fables concerning animal characters, actions, and morals that can be mapped to the elements in our frameworks, i.e., characters, actions, and proverbs. In terms of analogical mappings, they consider six categories of Shallow and Deep Attribute, Relational, Event, Structural, and Moral analogies. Moral analogies fall under system mappings in our framework, while others being more focused on the surface similarities can be mapped to specific mappings that we define based on characters, relationships, actions, goals, and locations — e.g. Event analogies to surface mappings based on actions — all of which fall under surface mappings. While certain scenarios might require further adaptations in the operationalization of the framework, we expect that the three-tier structure of the $ARN$ framework be used as the basis for both existing and future tasks, as the underlying concepts of surface/system mappings, far/near (dis)analogies, and the four partitions covering the interplay between them are well grounded in theory, and are well generalizable across domains and contexts.
@@ 4 $ARN$ Benchmark
In this section, we describe the realization of the framework in Section 3 as a benchmark for analogical reasoning over narratives ($ARN$). Without a loss of generality, we use narratives from the ePiC crowdsourced dataset (Ghosh and Srivastava, 2021) that consists of 2,500 short narratives (on average 4-8 sentences) annotated with 250 corresponding proverbs (10 narratives per proverb). We chose the ePiC dataset because of its systematic annotation of proverbs and the wide range of topics that it covers, allowing for the extraction of diverse elements with corresponding mappings.
In Section 4.1, we explain how we extract elements from the narratives in ePiC. The mappings between these elements are formed by quantifying their similarity across pairs of narratives (see Section 4.2). The formed surface mappings serve as a basis to find distractors (see Section 4.3), whereas the system mappings form corresponding analogies (see Section 4.4). The $ARN$ benchmark consists of ($Q,A,N$) triples, pairing analogies $A$ and their corresponding distractors $N$ for the same query narrative $Q$. The semantic similarity between $Q$ and $A$, and between $Q$ and $N$, are defined by using the number of their formed surface mappings. The statistics and quality analysis of $ARN$ are presented in Section 4.5. For more details about the involved preprocessing steps and examples, refer to Appendix A.

### v4

@@ 1 Introduction
Analogical reasoning is a core cognitive skill unique to humans (Penn et al., 2008; Hofstadter, 2001), defined as the ability to perceive and utilize the similarities between situations or events based on (systems of) relations rather than surface similarities (Holyoak, 2012; Gentner et al., 2012). The dichotomy between relational and surface similarity is illustrated in Figure 1: while the narratives $Q$ and $N$ overlap in terms of characters, locations, and actions (surface similarity), they fail to create a system of relational correspondences and thus are disanalogous as shown by their high-level message. Meanwhile, $Q$ and $A$ are dissimilar on the surface while forming a coherent relational system of correspondences through the high-level message: no pain, no gain. Mappings based on systems of relations (system mappings) have priority over independent relations (Gentner et al., 1993). Further, depending on whether the corresponding situations/events share the same domain or not, analogies and disanalogies can be qualified as near or far (Gentner, 1983).
@@
Analogy enables creative inferences, explanations, and generalization of knowledge and has been used for scientific inventions (Dunbar and Klahr, 2012), solving problems (Gick and Holyoak, 1980), and policy-making (Houghton, 1998). It has also been the subject of cognitive theories and studies about humans for common processes, such as retrieval of memories (Wharton et al., 1994) and problem-solving (Gick and Holyoak, 1980), mostly leveraging narratives as their experimental medium (e.g., Gentner and Toupin, 1986; Gentner et al., 1993; Wharton et al., 1994), given their multi-tiered nature and potential for abstraction. However, these studies focus on humans or models expecting well-structured inputs (Lu et al., 2022), are small-scale (Ichien et al., 2020, Table 5), and involve manually curated narratives, lacking an indication of how they can be applied to create scalable benchmarks for evaluating computational models. Nevertheless, borrowing from narratology theories Mani (2013); Gardner (2003), elements of narratives may enable analogical theories to be scaled and formalized as computational frameworks and corresponding benchmarks; however, this avenue has not been explored to date.
@@
Meanwhile, analogical reasoning has also been relatively popular in natural language processing (NLP), typically framed as an intelligence test for models compared against humans. So-called word-based, proportional analogies of the form $(A:B::C:D)$ (e.g., Mikolov et al., 2013a, b; Gladkova et al., 2016; Ushio et al., 2021) are often used to measure the potential of word embeddings and language models in terms of analogical reasoning. Recent studies (Webb et al., 2023) show a strong ability of state-of-the-art (SOTA) large language models (LLMs) to discover proportional word analogies, though this skill degrades with higher complexity (Wijesiriwardene et al., 2023) or associative phrasing of the input (Stevenson et al., 2023). Shifting toward more complex settings, narrative-based analogy benchmarks that involve system mappings rather than simple word-based relational mappings, and are aligned with cognitive theories have been rarely considered (Nagarajah et al., 2022; Wijesiriwardene et al., 2023), with limitations in scope and generalizability. Thus, we note a significant gap between the expressivity of cognitive theories of analogies and the present analogical reasoning benchmarks in NLP. Specifically, existing computational frameworks seldom include analogy over narratives and do not provide a formalization of key concepts such as system mappings and near/far analogs. Consequently, the benchmarks used to evaluate SOTA models do not provide the ability for systematic exploration of the sensitivity of model performance to different analogical categories beyond word-level relations.
@@ 2 Related Work
Many categorizations exist for analogies, with a common distinction between surface and relational (system) mappings (e.g., Halford, 1992; Gentner and Rattermann, 1991; Gentner, 1982; Premack, 1983). Gentner (1983)’s Structural Mapping Engine (SME) defines domains and situations as systems of objects, object attributes, and relations between objects, and distinguishes between mappings formed based on each. SME defines analogies as mappings based on relations rather than objects (attributes), prioritizing interconnected systems of mappings across domains (system mappings). Similarly, Holyoak and Thagard (1996) distinguish between three mapping categories of increasing complexity: attribute, relational, and system. Each mapping category can be formalized using a propositional representation with predicates (e.g., Mentor of) and filler subjects (e.g., Kim, Emily). Aligned with both frameworks, we categorize mappings into: surface mappings covering all categories of lower-order mappings (e.g., based on objects and object attributes), and system of relational mappings (system mappings) that create a coherent system of correspondences across two situations.
System mappings form analogies, having a higher priority over other mappings (Gentner, 1983; Holyoak and Thagard, 1996); yet, they can co-exist with surface mappings. Surface mappings decrease the cognitive load for recognizing system mappings since they create within-domain (near) analogies, also known as literal similarity (Gentner, 1983). Cross-domain (far) analogies, not forming surface mappings, are more challenging to identify (Alexieva and Hristova, 2017; Green et al., 2006, 2008, 2010), promote relational thinking (Vendetti et al., 2014), and are seen as more sound when making arguments (Gentner et al., 1993). Conversely, the sole existence of surface mappings yields near disanalogies or mere appearance, while the absence of both surface and system mappings is a far disanalogy, or dissimilarity (Gentner, 1983). While the distinction between near and far (dis)analogies is common in popular cognitive psychology (CogPsy) frameworks, it is unclear how to operationalize them at scale for narratives, a gap we bridge in this paper.
@@
To understand what may form surface and system mappings in narratives, we review frameworks that describe narratives through their key elements. Bal (2017) distill four narrative elements: actors, events, time, and location. Mani (2013) distinguish between characters, time, and plot. Building upon these categorizations, Vossen et al. (2021) expand the scope to include sequentiality and focalizer in addition to events and characters. More comprehensive and more widely adopted, Gardner (2003) consider characters, plot, theme, setting, point of view, and style. To our knowledge, no prior work has leveraged such elements to infer surface and system similarities across narratives. We base our categorization of elements on categories of Gardner (2003), given its comprehensiveness, to form surface and system mappings, inferring analogies and disanalogies across narratives.
@@
For both models and humans, evaluating analogical reasoning as an intelligence test (Hesse, 1965; Mitchell, 2021) is often posed as a word-based proportional task ($A:B::C:D$); having $Happy:Sad::C:D$, and asking to infer the missing $Angry:Calm$ or $Calm$ given $Happy:Sad::Angry:D$. Proportional analogy benchmarks (e.g., Turney et al., 2003; Jurgens et al., 2012; Mikolov et al., 2013a, b, c; Gao et al., 2014; Gladkova et al., 2016; Czinczoll et al., 2022; Kotchian and Simmons, 2012; Tutors, ) differ in scale (up to thousands of samples), number of relations (between one and few dozens), and domains (general domain or science). Depending on the context, SOTA LLMs perform worse than humans (Collier et al., 2023; Mitchell et al., 2023), or outperform humans (which might be artifacts of training data; Lewis and Mitchell, 2024) in word-based analogy tasks; however, they fail to perform well in story analogies (Webb et al., 2023). Notably, while cognitive theories of analogy often focus on narratives and study the distinction between surface, relational, and system similarities, existing work in NLP has a narrower focus as far as analogies between word pairs. In this study, we extend the focus from word-based relational analogies to relational system analogies in narratives.
@@
Meanwhile, focusing on the limited studies of matching, retrieving, or completing more complex analogies in narratives in CogPsy (Ichien et al., 2020): Wharton et al. (1994) develop 14 sets of four stories with different themes and events, used to compare human analogies between themes and events; Wijesiriwardene et al. (2023) study analogies in stories using negation, entailment, and explanations, using available datasets for each; Gentner and Toupin (1986) study the effect of systematicity and surface similarity of narratives’ characters on analogical reasoning with 54 narratives; Gick and Holyoak (1980) study analogical reasoning in problem-solving and assess the effect of similarities in problems and solutions separately. These tasks consider narrative analogies formed by humans, are thus limited in size (dozens of samples), and do not exploit the richness of the frameworks on narrative elements. Although large-scale story corpora exist (e.g., Mostafazadeh et al., 2016; Storks et al., 2021; Andrus et al., 2022), analogies are extremely scarce in free texts (Sultan and Shahaf, 2022), and these resources are not adequate for evaluating analogical reasoning (Wijesiriwardene et al., 2023) since they do not consider the theoretical dimensions of analogical reasoning from a CogPsy perspective, which is a gap we cover in this study. Most similar to our work, Nagarajah et al. (2022) formalize six cognitively-inspired dimensions of analogical reasoning over hundreds of Aesop’s fables. However, while these dimensions can be mapped to our framework, their fine granularity limits their application to the set of fables they are devised for. Instead, we formulate a comprehensive framework that formalizes analogies and disanalogies by using narrative elements from narratology. This framework explicitly connects to the CogPsy theoretical dimensions of analogical reasoning frameworks, resulting in the inception of the first extensive, theory-grounded benchmark for analogical reasoning within narratives. Concurrently to our work, Jiayang et al. (2023) focus on analogies in narratives by providing a large-scale benchmark built on top of existing resources; however, they focus specifically on the distinction between surface and relational similarities. We go one step further by covering system mappings and the possible interactions between near/far analogies and distractors to paint a more accurate picture of models’ analogical reasoning abilities.
@@ 3 $ARN$ Framework
Inspired by analogical reasoning frameworks in CogPsy (e.g., Wharton et al., 1994; Gentner et al., 1993), we formulate analogical reasoning over narratives as a task where input consists of a query narrative $Q$ and two candidate narratives, $A$ and $N$ (see Figure 1). The goal is to select the narrative $A$ that is analogous to $Q$ rather than the disanalogous one $N$, based on system mappings. We frame it as a binary classification task to enable an unambiguous evaluation of the analogical reasoning abilities of models and humans (Frank, 2023). We realize this task through a compositional framework with three components (Figure 2): 1. extraction of elements of narratives, such as characters, actions, and high-level messages (see Section 3.1); 2. formation of surface and system mappings based on those elements (see Section 3.2); 3. task formulation, by defining (dis)analogies based on the mappings and partitioning the data points according to their semantic distance to the query narratives (see Section 3.3).
@@ 3.1 Elements of Narratives
We take the proverbial or high-level message of a narrative as an instantiation of its overarching theme (Gardner, 2003), which encapsulates individual elements of narratives into a coherent idea that the writer wants to convey. More formally, proverbs are short sentences containing wisdom, truth, morals, and traditional views in a metaphorical form (Mieder, 1993). For instance, query narrative $Q$ in Figure 1 is associated with the proverb: no pain, no gain.
@@ 3.2 Mappings between Narratives
Unlike surface mappings that focus on similarities between independent elements, system mappings are formed by interconnected correspondences of relations across two narratives (Gentner, 1983). This system of correspondences can be characterized and formed by leveraging narratives’ proverbs (Wijesiriwardene et al., 2023); hence, two narratives with the same proverb are connected by system mappings. Similar to the moral dimensions of Nagarajah et al. (2022), proverbs provide an abstract summary of the causal structure of a narrative, acting as mediators for the transition between two analogs.
@@ 3.4 Generalizability of ARN to Narrative Analogies
The purpose of $ARN$ is to provide a generalizable framework to study analogical reasoning in narratives grounded in narratology and analogical reasoning in CogPsy research. The provided set of elements is universal in narratives (Gardner, 2003): entities’ attributes or surface relations, proverbs as overarching messages of narratives, together with their corresponding surface and system mappings, can be broadly found in any narrative. This universality makes $ARN$ an adequate framework to describe other existing datasets. To illustrate this point, we describe how two existing benchmarks for narrative analogies can be mapped to the elements of $ARN$. StoryAnalogy (Jiayang et al., 2023) uses characters, actions, and relationships of narratives to represent and distinguish between entity and relational similarities while not covering higher-order mappings of analogical messages. The entity and relational elements of StoryAnalogy directly correspond to the elements in $ARN$, whereas higher-order elements and mappings, such as proverbs, are not present in this dataset. Similarly, Nagarajah et al. (2022) capture structural similarities between Aesop’s fables concerning animal characters, actions, and morals that can be mapped to the elements in our frameworks, i.e., characters, actions, and proverbs. In terms of analogical mappings, they consider six categories of Shallow and Deep Attribute, Relational, Event, Structural, and Moral analogies. Moral analogies fall under system mappings in our framework, while others, being more focused on the surface similarities, can be mapped to specific mappings that we define based on characters, relationships, actions, goals, and locations — e.g., Event analogies to surface mappings based on actions — all of which fall under surface mappings. While certain scenarios might require further adaptations in the operationalization of the framework, we expect that the three-tier structure of the $ARN$ framework to be used as the basis for both existing and future tasks, as the underlying concepts of surface/system mappings, far/near (dis)analogies, and the four partitions covering the interplay between them are well grounded in theory and generalizable across domains and contexts.
@@ 4 $ARN$ Benchmark
In this section, we describe the realization of the framework in Section 3 as a benchmark for Analogical Reasoning over Narratives ($ARN$). Without loss of generality, we use narratives from the ePiC crowdsourced dataset (Ghosh and Srivastava, 2022) that consists of 2,500 short narratives (on average 4-8 sentences) annotated with 250 corresponding proverbs (10 narratives per proverb). We chose the ePiC dataset because of its systematic annotation of proverbs and the wide range of topics that it covers, allowing for the extraction of diverse elements with corresponding mappings.
In Section 4.1, we explain how we extract elements from the narratives in ePiC. The mappings between these elements are formed by quantifying their similarity across pairs of narratives (see Section 4.2). The formed surface mappings serve as a basis to find distractors (see Section 4.3), whereas the system mappings form corresponding analogies (see Section 4.4). The $ARN$ benchmark consists of ($Q,A,N$) triples, pairing analogies $A$ and their corresponding distractors $N$ for the same query narrative $Q$. The semantic similarity between $Q$ and $A$, and between $Q$ and $N$, are defined by using the number of formed surface mappings. The statistics and quality analysis of $ARN$ are presented in Section 4.5. For more details about the involved preprocessing steps and examples, refer to Appendix A.